Once ignited, a small rocket motor on a spacecraft exerts a constant force of 10 N for 7.80 s. During the burn the rocket causes the 100-kg craft to accelerate uniformly. Determine that acceleration
1
Expert's answer
2020-07-20T14:56:46-0400

We can determine the acceleration of the rocket from the Newton's Second Law of Motion:


"F=ma,""a=\dfrac{F}{m} = \dfrac{10 \ N}{100 \ kg} = 0.1 \ \dfrac{m}{s^2}."

Answer:

"a=0.1 \ \dfrac{m}{s^2}."